OpenStudy (anonymous):

solve for n m/9 + n/8 = 5

OpenStudy (anonymous):

n/8=5-m/9 n/8=(45-m)/9 n=8(45-m)/9 n=(360-8m)/9

Directrix (directrix):

n = 40 - (8/9) m which is equivalent to n=(360-8m)/9 from vengeance921